The empirical formula of the compound is CH2O . Its molecular weight is 180.The molecular formula of the compound is

Grade:12

3 Answers

pratham ashish
14 years ago

Hi,

 

the wt of emperical = 12+2+ 16= 30

molecular formula = n ( emperical formula)

 

the way to calculate n:

n =   molecular wt of the compound /  wt of emperical formula

 

here we hav  n = 180/30 = 6

hence the molecular formula = 6 ( emperical formula)

i.e  6(CH2O) = C6H12O6

 

good luck & enjoy ur preparation

Ramesh V
14 years ago

atomic weight of C=12 ,H=1,O=16 ,

mol. wt. of CH2O is 30

given mol. wt. of compound (nCH2O) is 180,

n=180/30=6

so, compound is C6H12O6

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh  
IIT Kgp - 05 batch




Pratham Ashish
14 years ago

hi,

 empirical wt . of compound = 12 + 2 + 16

                                         = 30

 mol. wt is 180 

thw ratio between the two wt. is called n

          n = 180/ 30

             =6 

so, the mol. formula of compoundwill be 

                                           =  ( CH2 o ) 6

                                            = C6 H12 O6